# Flexible Tool Selection through Low-dimensional Attribute Alignment of Vision and Language

### v2

@@ Abstract
Flexible tool selection reflects a complex cognitive ability that distinguishes humans from other species, yet computational models that capture this ability remain underdeveloped. We developed a framework using low-dimensional attribute representations to bridge visual tool perception and linguistic task understanding. We constructed a comprehensive dataset (ToolNet) containing 115 common tools labeled with 13 carefully designed attributes spanning physical, functional, and psychological properties, paired with natural language scenarios describing tool usage. Visual encoders (ResNet/ViT) extract attributes from tool images while fine-tuned language models (GPT-2, LLaMA, DeepSeek) derive required attributes from task descriptions. Our approach achieves 74% accuracy in tool selection tasks—significantly outperforming direct tool matching (20%) and smaller multimodal models (21%-58%), while approaching performance of much larger models like GPT-4o (73%) with substantially fewer parameters. Ablation studies revealed that manipulation-related attributes (graspability, hand-relatedness, elongation) consistently prove most critical across modalities. This work provides a parameter-efficient, interpretable solution that mimics human-like tool cognition, advancing both cognitive science understanding and practical applications in tool selection tasks.
@@ I Introduction
The ability to flexibly select and use tools represents a remarkable cognitive capability that extends human physical limitations and sets us apart from other species[1, 2]. While certain non-human animals demonstrate rudimentary tool usage—such as chimpanzees inserting sticks into termite mounds[3], orangutans using long poles to retrieve fruit[4], or ants employing leaves to transport food[5]-humans exhibit a uniquely flexible and sophisticated capacity for tool manipulation across diverse contexts.
Unlike animals, humans can design complex multi-component tools[6], transmit tool-making knowledge across generations through language[7], adapt tools for purposes far removed from their original function[8, 9], and create abstract tools like mathematical symbols and computer algorithms[10]. This flexibility allows humans to select appropriate tools for novel situations, repurpose objects for unintended functions, and even create new tools to address emerging challenges. Despite its evolutionary significance and centrality to human cognition, the computational and neural mechanisms underlying flexible tool selection remain insufficiently understood[11, 12, 13, 14].
@@
Despite these insights from cognitive neuroscience, computational models that effectively capture this attribute-based flexible tool selection mechanism remain underdeveloped[18]. Current approaches to modeling tool selection often rely on either direct mapping between task descriptions and tool labels[19] or comprehensive multimodal processing that requires extensive computational resources. Additionally, the absence of standardized datasets connecting tool images, usage scenarios, and underlying attributes has impeded progress in this domain.
@@
    We introduce a carefully designed 13-dimensional attribute space that captures both physical properties (elongation, size, hardness) and functional characteristics (graspability, body extension) of tools. We construct comprehensive datasets to support attribute-based tool selection research: (1) a tool image-attribute dataset containing 115 common tools with 13 corresponding attribute ratings, (2) a tool scenario-attribute dataset featuring textual descriptions of tool usage scenarios paired with corresponding attribute requirements, and (3) a tool matching test set comprising 100 scenario descriptions with 10 candidate tool images each.
@@
    We demonstrate that our attribute-based approach achieves 74% accuracy in tool selection tasks, substantially outperforming direct tool name matching (20%) and smaller multimodal large language models (LLMs) (21%-58%), while showing competitive performance against much larger multimodal LLMs like GPT-4o[21] (73%) and Gemini-2.0-Pro[22] (72%), despite using significantly fewer parameters.
This work provides a computationally efficient and cognitively plausible approach to flexible tool selection. By decomposing tool selection into attribute-based representations significantly enhances model performance while requiring substantially fewer parameters than large-scale multimodal LLMs. Through ablation studies, we identify key attributes driving model performance, with functional properties like graspability, elongation, and hand-relatedness proving most critical for accurate tool selection. This work bridges cognitive science and artificial intelligence (AI) by implementing a neurally-inspired computational framework for flexible tool use. By demonstrating the efficacy of attribute-based representations in both visual and linguistic domains, our research provides insights into potential mechanisms underlying human tool cognition while offering a practical system that can directly analyze any scenario, process images of available tools, and select the most appropriate tool for the given context.
@@ II-C Transfer Learning and Task Alignment
In computer vision, transfer learning has become the dominant paradigm for downstream task adaptation [51, 52]. Convolutional architectures like ResNet [53] demonstrated remarkable transferability of learned features across diverse visual tasks, while Vision Transformers (ViT)[54] further improved this capability through their attention-based feature extraction. For adapting these pretrained visual backbones to specific downstream objectives, several approaches have proven effective. Linear probing trains only the final classification layer while keeping backbone weights fixed [55], offering computational efficiency with minimal risk of catastrophic forgetting [56]. Full fine-tuning adjusts all parameters but typically applies lower learning rates to pretrained layers, enabling more comprehensive adaptation when sufficient task data is available [57]. More parameter-efficient approaches include adapter modules [58] that insert small trainable components between frozen layers, and knowledge distillation techniques [59] that transfer capabilities from larger teacher models to more efficient student networks.
@@ III-A Dataset Construction
The effectiveness of our framework relies heavily on a well-designed attribute space and comprehensive datasets. Our 13-dimensional attribute space was carefully designed to capture the essential physical, functional, and psychological properties of tools. These dimensions were selected based on their theoretical relevance to tool cognition and empirical evidence of their importance in human tool selection processes. The attributes can be broadly categorized into three groups: physical properties (elongation, spiky, size, smoothness, texturedness, and hardness) that characterize the intrinsic material and structural characteristics of tools; functional properties (graspability, hand involvement, force requirements, and body extension) that describe how the tool interfaces with human users; and psychological properties (threatness, valence, and arousal) that represent the emotional and psychological aspects of tool interaction.
@@
This multi-faceted dataset design enables us to train models to extract attributes from both visual tool representations and linguistic scenario descriptions, while also providing a robust benchmark for evaluating cross-modal tool selection performance.
@@ III-B Problem Formulation
$s_{cos}(\mathbf{a}_{d},\mathbf{a}_{t})=\frac{\mathbf{a}_{d}\cdot\mathbf{a}_{t}}{||\mathbf{a}_{d}||\cdot||\mathbf{a}_{t}||}$ (3)
$s_{euc}(\mathbf{a}_{d},\mathbf{a}_{t})=-||\mathbf{a}_{d}-\mathbf{a}_{t}||_{2}$ (4)
@@
III-C Vision-Language Model
@@
III-C1 Visual Encoder
The visual encoder follows a two-stage architecture comprising a pre-trained feature extractor backbone followed by an attribute prediction head, corresponding to the visual pathway in Fig.1(b). We experiment with three backbone architectures: ResNet-18, ResNet-50, and Vision Transformer (ViT-B/16), all initialized with ImageNet pre-trained weights. This design enables efficient transfer learning while maintaining the model’s capacity to extract tool-specific attributes.
The attribute prediction head is implemented as a multi-layer perceptron (MLP) that maps high-dimensional visual features (512/2048/768-dimensional, depending on the backbone) to our 13-dimensional attribute space. The MLP architecture consists of three fully connected layers with dimensions 256 and 64 for the hidden layers, followed by layer normalization and ReLU activation functions after each hidden layer except the final output layer. This transformation allows the model to distill relevant functional and physical properties from complex visual representations.
III-C2 Language Encoder
@@ IV Results
Figure 3: Performance evaluation of visual and language models. (a) Test accuracy of visual models (ResNet18, ResNet50, ViT-B/16) on attribute prediction and most similar class identification tasks. ResNet50 achieves the highest performance with 96.05% attribute-wise accuracy and 92.70% most similar classes accuracy. (b) Training and testing attribute-wise accuracy of language models (GPT-2, LLaMA-3.2-1.2B, DeepSeek-R1-1.5B) across different dataset sizes. The small dataset contains 10 training scenarios and 3 testing scenarios per tool, the middle dataset contains 90 training scenarios and 10 testing scenarios per tool, and the large dataset contains 475 training scenarios and 25 testing scenarios per tool. ”GPT-2 with quiz” represents adding the question ”What tool is relevant to this scene?” to scenario descriptions, which decreased performance compared to standard GPT-2. Larger models and datasets yield better performance, with DeepSeek-R1-1.5B demonstrates the best performance with 74.34% attribute-wise accuracy on the largest dataset.
@@ IV-A Visual Model Performance
We evaluated performance using two complementary metrics. First, attribute-wise accuracy measures the model’s ability to predict individual attribute values accurately on the 7-point scale. Specifically, the predicted values are rounded to the nearest integer, and the prediction is considered correct only if it exactly matches the ground truth value. ResNet50 achieved 96.05% accuracy across all attributes and test samples, outperforming both ResNet18 (93.01%) and ViT-B/16 (94.34%). Second, most similar class accuracy evaluates whether the predicted attribute vector’s closest matching tool category (by cosine similarity) matches the ground truth category. ResNet50 again demonstrated superior performance (92.70%), closely followed by ViT-B/16 (92.61%) and ResNet18 (89.40%).
@@ IV-B Language Model Performance
Figure 5: Ablation study on attributes for tool selection. This experiment examines how removing individual attributes affects performance in both visual models, language models, and the combined system. Results show that functional attributes (particularly graspability, elongation, and hand-relatedness) have the greatest impact on individual model performance, while their removal can improve the combined model’s accuracy. In contrast, attributes like valence, spiky, size, body extension, and arousal have minimal impact on model performance, suggesting certain attributes are more critical than others for flexible tool selection.
@@ IV-D Ablation Studies
To understand the relative importance of different attributes in our framework, we conducted an ablation study where we systematically removed individual attributes and measured the impact on model performance. Fig. 5 presents these results for both individual encoders and the combined tool selection task.
@@
For the visual encoder, consistent patterns emerged across all architectures. Hand-relatedness removal caused the most substantial performance degradation (12.35-13.39%), followed closely by elongation (11.91-12.61%) and graspability (9.30-10.35%). In contrast, attributes like valence, size, and spikiness showed minimal impact when removed (generally below 3% decrease). This suggests that functional characteristics related to human interaction and shape properties (particularly elongation) are the most informative visual cues for distinguishing between tool categories.
@@
Similar patterns appeared in language models, where elongation, hand-relatedness, and graspability consistently proved most critical for accurate prediction. For DeepSeek-R1-1.5B, removing elongation decreased accuracy by 7.10%, while hand-relatedness and graspability reduced accuracy by 6.23% and 5.67%, respectively. This consistency across modalities suggests that functional attributes related to tool manipulation and physical form are inherently more distinguishing in both visual and linguistic representations.
In the end-to-end tool selection task, an interesting pattern emerged. While most attribute removals resulted in slight performance decreases (0-3%), removing hand-relatedness unexpectedly improved accuracy by 4%. This suggests potential interactions between attributes in cross-modal matching that differ from their contributions in isolated visual or language pathways. The small impact of individual attribute removal on end-to-end performance indicates redundancy in the attribute representation, allowing the system to maintain reasonable performance even with missing dimensions.
@@
These ablation studies reveal that our attribute space effectively captures the most salient properties for flexible tool selection, with functional and manipulation-related attributes proving particularly critical across modalities.

### v4

@@ Abstract
Flexible tool selection reflects a complex cognitive ability that distinguishes humans from other species, yet computational models that capture this ability remain underdeveloped. We developed a framework using low-dimensional attribute representations to bridge visual tool perception and linguistic task understanding. We constructed a dataset (ToolNet) containing 115 common tools labeled with 13 carefully designed attributes spanning physical, functional, and psychological properties, paired with natural language scenarios describing tool usage. Visual encoders (ResNet/ViT) extract tool attributes from images while fine-tuned language models (GPT-2, LLaMA, DeepSeek) derive required attributes from task descriptions. Our approach achieves 74% accuracy in tool selection tasks—significantly outperforming simpler baselines (20%) and smaller multimodal models (21%-58%), while approaching performance of much larger models like GPT-4o (73%) with substantially fewer parameters. Human evaluation studies validate our framework’s alignment with human decision-making patterns, and generalization experiments demonstrate effective performance on novel tool categories. Ablation studies revealed that manipulation-related attributes (graspability, elongation, hand-relatedness) consistently prove most critical across modalities. This work provides a parameter-efficient, interpretable solution that mimics human-like tool cognition, advancing both cognitive science understanding and practical applications in tool selection tasks.
@@ I Introduction
The ability to flexibly select and use tools represents a remarkable cognitive capability that extends human physical limitations and sets us apart from other species[1, 2]. While certain non-human animals demonstrate rudimentary tool usage—such as chimpanzees inserting sticks into termite mounds[3], orangutans using long poles to retrieve fruit[4], or ants employing leaves to transport food[5]. This fundamental difference in cognitive flexibility represents a key evolutionary advancement that enables humans to exhibit uniquely sophisticated capacity for tool manipulation across diverse contexts.
Beyond basic tool use, humans demonstrate several unique capabilities: designing complex multi-component tools[6], transmitting tool-making knowledge across generations through language[7], adapting tools for purposes far removed from their original function[8, 9], and creating abstract tools like mathematical symbols and computer algorithms[10]. This flexibility allows humans to select appropriate tools for novel situations, repurpose objects for unintended functions, and create new tools to address emerging challenges. Despite its evolutionary significance and centrality to human cognition, the computational and neural mechanisms underlying flexible tool selection remain poorly understood[11, 12, 13, 14].
@@
Despite these insights from cognitive neuroscience, computational models that effectively capture this attribute-based flexible tool selection mechanism remain underdeveloped[18]. Current approaches to modeling tool selection often rely on either direct mapping between task descriptions and tool labels[19] or multimodal processing that requires extensive computational resources. Additionally, the absence of standardized datasets connecting tool images, usage scenarios, and underlying attributes has impeded progress in this domain.
@@
    We introduce a carefully designed 13-dimensional attribute space that captures both physical properties (elongation, size, hardness) and functional characteristics (graspability, body extension) of tools. We construct four datasets to support attribute-based tool selection research: (1) a tool image-attribute dataset containing 115 common tools with 13 corresponding attribute ratings, (2) a tool scenario-attribute dataset featuring textual descriptions of tool usage scenarios paired with corresponding attribute requirements, (3) a tool matching test set comprising 100 scenario descriptions with 10 candidate tool images each, complemented by human evaluation data from 30 participants, and (4) a novel tool generalization dataset featuring 25 previously unseen tool categories for testing cross-category adaptability.
@@
    We demonstrate that our attribute-based approach achieves 74% accuracy in tool selection tasks, substantially outperforming direct tool name matching (20%) and smaller multimodal large language models (LLMs) (21%-58%), while showing competitive performance against much larger multimodal LLMs like GPT-4o[21] (73%) and Gemini-2.0-Pro[22] (72%), despite using significantly fewer parameters. Human evaluation studies confirm alignment between our model predictions and human decision-making patterns. Additionally, generalization experiments on novel tool categories demonstrate the framework’s ability to extend beyond training data through attribute-based reasoning.”
Generalization experiments on novel tool categories further demonstrate that our attribute-based framework can effectively reason about previously unseen tools. This work provides a computationally efficient and cognitively plausible approach to flexible tool selection. By decomposing tool selection into attribute-based representations significantly enhances model performance while requiring substantially fewer parameters than large-scale multimodal LLMs. Through ablation studies, we identify key attributes driving model performance, with functional properties like graspability, elongation, and hand-relatedness proving most critical for accurate tool selection. By demonstrating the efficacy of attribute-based representations in both visual and linguistic domains, our research provides insights into potential mechanisms underlying human tool cognition while offering a practical system that can directly analyze any scenario, process images of available tools, and select the most appropriate tool for the given context.
@@ II-C Transfer Learning and Task Alignment
In computer vision, transfer learning has become the dominant paradigm for downstream task adaptation [51, 52]. Convolutional architectures like ResNet [53] demonstrated remarkable transferability of learned features across diverse visual tasks, while Vision Transformers (ViT)[54] further improved this capability through their attention-based feature extraction. For adapting these pretrained visual backbones to specific downstream objectives, several approaches have proven effective. Linear probing trains only the final classification layer while keeping backbone weights fixed [55], offering computational efficiency with minimal risk of catastrophic forgetting [56]. Full fine-tuning adjusts all parameters but typically applies lower learning rates to pretrained layers, enabling more effective adaptation when sufficient task data is available [57]. More parameter-efficient approaches include adapter modules [58] that insert small trainable components between frozen layers, and knowledge distillation techniques [59] that transfer capabilities from larger teacher models to more efficient student networks.
@@ III-A Dataset Construction
To enable effective tool selection, we first identified the requirements for bridging visual tool perception and linguistic task understanding. Based on these requirements, we designed a 13-dimensional attribute space that captures the essential physical, functional, and psychological properties of tools. These dimensions were selected based on their theoretical relevance to tool cognition and empirical evidence of their importance in human tool selection processes. The attributes can be broadly categorized into three groups: physical properties (elongation, spikiness, size, smoothness, texturedness, and hardness) that characterize the intrinsic material and structural characteristics of tools; functional properties (graspability, hand involvement, force requirements, and body extension) that describe how the tool interfaces with human users; and psychological properties (threatness, valence, and arousal) that represent the emotional and psychological aspects of tool interaction.
@@
To establish human performance baselines and evaluate model-human alignment, we conducted a human evaluation study with 30 participants. Each participant evaluated all 100 scenarios from the Tool Matching Dataset, rating the appropriateness of each of the 10 candidate tools for the given usage scenario. Participants used a 7-point Likert scale, where a rating of 7 indicates that the tool is highly suitable for the described scenario, while a rating of 1 indicates complete unsuitability. This evaluation protocol generated 30,000 individual ratings (30 participants × 100 scenarios × 10 tools), providing rich data for analyzing both human decision-making patterns and inter-rater reliability. The study received approval from the institutional ethics review board, and all participants provided informed consent before participation.
To assess our framework’s ability to generalize to previously unseen tools, we constructed a New Tool Matching Dataset featuring 25 completely new tool categories not present in the original ToolNet collection. These novel tools, illustrated in Fig. S5, were selected to cover diverse functional categories. For each novel tool, we generated 20 unique usage scenarios using the same protocol employed for the original Tool Scenario-Attribute Dataset, resulting in 500 test scenarios total (25 tools × 20 scenarios). Each test case in the New Tool Matching Dataset consists of a scenario description involving the novel tool paired with 10 candidate tool images: 1 target image of the novel tool and 9 distractor images randomly selected from the original ToolNet collection. This design enables evaluation of whether our attribute-based framework can correctly identify novel tools based on scenario descriptions, even when the visual encoder has never encountered these specific tool categories during training.
This multi-faceted evaluation approach enables comprehensive assessment of our framework across three critical dimensions: computational performance on standard benchmarks, alignment with human decision-making processes, and generalization capability to novel tool categories.
@@ III-B Problem Formulation
[…]
@@
III-C Correlation Analysis
To evaluate the consistency between human evaluators and the similarity between model predictions and human judgments, we employ two complementary correlation measures.
For assessing inter-rater reliability among human evaluators, we calculate Pearson correlation coefficients between all pairs of the 30 human participants. Given two raters $i$ and $j$ with rating vectors $\mathbf{r}_{i}=[r_{i,1},r_{i,2},...,r_{i,n}]$ and $\mathbf{r}_{j}=[r_{j,1},r_{j,2},...,r_{j,n}]$ across $n$ tool-scenario pairs, the Pearson correlation coefficient is computed as:
$\rho_{ij}=\frac{\sum_{k=1}^{n}(r_{i,k}-\bar{r}_{i})(r_{j,k}-\bar{r}_{j})}{\sqrt{\sum_{k=1}^{n}(r_{i,k}-\bar{r}_{i})^{2}}\sqrt{\sum_{k=1}^{n}(r_{j,k}-\bar{r}_{j})^{2}}}$ (6)
where $\bar{r}_{i}$ and $\bar{r}_{j}$ are the mean ratings for raters $i$ and $j$, respectively. We compute pairwise correlations for all $\binom{30}{2}=435$ rater pairs and analyze the distribution of correlation coefficients to assess overall rating consistency.
To evaluate the similarity between model predictions and human preferences, we employ Spearman rank correlation, which measures the monotonic relationship between rankings regardless of the specific numerical values. For each scenario, we obtain both human ratings (averaged across all 30 participants) and model-predicted similarity scores for the 10 candidate tools. Let $\mathbf{h}=[h_{1},h_{2},...,h_{10}]$ represent the average human ratings and $\mathbf{m}=[m_{1},m_{2},...,m_{10}]$ represent the model scores for the 10 tools in a given scenario. The Spearman correlation coefficient is calculated as:
$\rho_{s}=1-\frac{6\sum_{k=1}^{10}d_{k}^{2}}{10(10^{2}-1)}$ (7)
where $d_{k}=\text{rank}(h_{k})-\text{rank}(m_{k})$ is the difference between the ranks of tool $k$ in the human ratings versus model predictions. We compute the Spearman correlation for each scenario and report the mean correlation across all scenarios to quantify overall alignment between model and human tool preferences.
III-D Vision-Language Model
@@
III-D1 Visual Encoder
The visual encoder follows a two-stage architecture comprising a pre-trained feature extractor backbone followed by an attribute prediction head, corresponding to the visual pathway in Fig.1(b). We experiment with three backbone architectures: ResNet-18, ResNet-50, and Vision Transformer (ViT-B/16), all initialized with ImageNet pre-trained weights. This design enables efficient transfer learning while preserving the model’s capacity to extract tool-specific attributes.
The attribute prediction head is implemented as a multi-layer perceptron (MLP) that maps high-dimensional visual features (512/2048/768-dimensional, depending on the backbone) to our 13-dimensional attribute space. The MLP architecture consists of three fully connected layers: input features → 256 → 64 → 13 (attribute dimensions), with layer normalization and ReLU activation functions after each hidden layer. This transformation allows the model to distill relevant functional and physical properties from complex visual representations.
III-D2 Language Encoder
@@ IV Results
Figure 3: Performance evaluation of visual and language models. (a) Test accuracy of visual models (ResNet18, ResNet50, ViT-B/16) on attribute prediction and most similar class identification tasks. ResNet50 achieves the highest performance with 96.05% attribute-wise accuracy and 92.70% most similar classes accuracy. (b) Training and testing attribute-wise accuracy of language models (GPT-2, LLaMA-3.2-1.2B, DeepSeek-R1-1.5B) across different dataset sizes. The small dataset contains 10 training scenarios and 3 testing scenarios per tool, the medium dataset contains 90 training scenarios and 10 testing scenarios per tool, and the large dataset contains 475 training scenarios and 25 testing scenarios per tool. ”GPT-2 with quiz” represents adding the question ”What tool is relevant to this scene?” to scenario descriptions, which decreased performance compared to standard GPT-2. Larger models and datasets yield better performance, with DeepSeek-R1-1.5B demonstrates the best performance with 74.34% attribute-wise accuracy on the largest dataset.
@@ IV-A Visual Model Performance
We evaluated performance using two complementary metrics. First, attribute-wise accuracy measures the model’s ability to predict individual attribute values accurately on the 7-point scale. Specifically, the predicted values are rounded to the nearest integer on the 7-point scale, and the prediction is considered correct only if it exactly matches the ground truth value. ResNet50 achieved 96.05% accuracy across all attributes and test samples, outperforming both ResNet18 (93.01%) and ViT-B/16 (94.34%). Second, most similar class accuracy evaluates whether the predicted attribute vector’s closest matching tool category (by cosine similarity) matches the ground truth category. ResNet50 again demonstrated superior performance (92.70%), closely followed by ViT-B/16 (92.61%) and ResNet18 (89.40%).
@@ IV-B Language Model Performance
Figure 5: Evaluation of attribute importance, human performance, and model generalization. (a) Attribute ablation study showing the impact of removing individual attributes on performance across visual models, language models, and the combined system. Functional attributes (graspability, elongation, hand-relatedness) show the greatest impact. (b) Human performance evaluation (n=30) on 100 scenarios. Score distributions reveal scenarios with clear tool preferences (orange) versus comparable alternatives (green). Inset: distribution of inter-rater correlations (mean r = 0.59). (c) Comparison of mean Spearman correlation with human ratings across different models, analyzed specifically for scenarios with small preference gaps (separation $\leq 1.8$). Our deepseek_resnet50_cosine approach (r = 0.373) outperforms smaller models (qwen2_5_vl_7b_instruct and deepseek-r1-1.5b) but shows lower correlation than large multimodal models (Gemini_2_0_Pro and GPT_4o). (d) Generalization to 25 novel tools using our deepseek_resnet50_cosine approach shows variable success rates across individual tool categories. Inset: Top-k accuracy comparison across different models.
@@ IV-D Ablation Studies
To understand the relative importance of different attributes in our framework, we conducted an ablation study on an extended dataset of 1,000 tool selection scenarios, systematically removing individual attributes and measuring the impact on model performance. Fig. 5(a) presents these results for visual encoders, language encoders, and the combined tool selection task.
@@
For the visual encoder, consistent patterns emerged across all architectures. Hand-relatedness removal caused the most substantial performance degradation (12.35-13.39%), followed closely by elongation (11.91-12.61%) and graspability (9.30-10.35%). In contrast, attributes like valence, spikiness, and size showed minimal impact when removed (generally below 3% decrease). This suggests that functional characteristics related to human interaction and shape properties (particularly elongation) are the most informative visual cues for distinguishing between tool categories.
@@
Similar patterns appeared in language models, where graspability, elongation, and hand-relatedness consistently proved most critical for accurate prediction. For DeepSeek-R1-1.5B, removing these functional attributes led to performance drops of 5-8%, while psychological attributes like valence and arousal showed minimal impact. This consistency across modalities suggests that functional attributes related to tool manipulation and physical form are inherently more distinguishing in both visual and linguistic representations.
@@
Consistent with the findings from the individual visual and language encoders, the ablation results for the end-to-end tool selection task reinforce the same attribute importance hierarchy. The removal of functional attributes resulted in the most significant performance degradation, while the exclusion of psychological attributes had a negligible impact. This reveals that our attribute space effectively captures the most salient properties for flexible tool selection, with functional and manipulation-related attributes proving particularly critical across modalities.
IV-E Model-Human Alignment
To evaluate our model’s alignment with human decision-making, we conducted a evaluation with 30 human participants on the Tool Matching Dataset. Fig. 5(b) presents the results of this human evaluation study. Human participants achieved 63% accuracy in selecting the most appropriate tool from 10 candidates. The score distribution analysis reveals two distinct patterns: scenarios where participants showed comparable preferences for the top-ranked and second-ranked tools (separation $\leq 1.8$, shown in green), indicating multiple viable tool options, and scenarios with clear preference gaps (separation >1.8>1.8, shown in orange), where one tool was decisively preferred over alternatives. Inter-rater reliability analysis showed moderate but significant agreement among human evaluators, with pairwise Pearson correlations averaging r = 0.59 ($p<0.001$).
To assess model-human alignment, we calculated Spearman rank correlations between model predictions and averaged human ratings. Notably, the correlation analysis in Fig. 5(c) focuses specifically on scenarios with small preference gaps (separation $\leq 1.8$), where multiple tools were considered viable alternatives by human participants. Fig. 5(c) shows that our attribute-based approach (deepseek_resnet50_cosine) achieved a mean correlation of r = 0.373 with human preferences. While this correlation is lower than large multimodal models like GPT-4o and Gemini-2.0-Pro, it substantially outperforms smaller baseline models, including the direct language approach (deepseek-r1-1.5b) and smaller multimodal models (qwen2_5_vl_7b_instruct).
IV-F Generalization to Novel Tools
To evaluate our framework’s ability to generalize beyond the training tool categories, we constructed a New Tool Matching Dataset featuring 25 completely new tool categories (Fig. S5). This evaluation tests whether our attribute-based approach can correctly identify appropriate tools for described scenarios even when encountering tool categories never seen during training.
Fig. 5(d) presents the success rates for individual novel tools, revealing substantial variation in generalization performance. Some tools achieved success rates well above the 10% chance level, with certain categories like cello and unicycle showing particularly strong performance. However, other novel tools, particularly those with unique functional properties or specialized applications, achieved success rates below chance level, indicating challenges in generalizing to tools with attribute combinations not well-represented in the training data.
Overall performance metrics demonstrate reasonable but limited generalization capability compared to large multimodal models. Our framework achieved 22.6% Top-1 accuracy, 36.8% Top-2 accuracy, and 51.0% Top-3 accuracy on the novel tool dataset. In comparison, Qwen2.5-VL-7B-Instruct achieved 21.0% Top-1 accuracy, 28.0% Top-2 accuracy, and 29.0% Top-3 accuracy, while Gemini-2.0-Pro demonstrated superior performance with 91.2% Top-1 accuracy, 97.6% Top-2 accuracy, and 98.8% Top-3 accuracy. While our results exceed both chance performance (10% for Top-1) and the smaller multimodal baseline, they indicate significant room for improvement in generalizing to completely unseen tool categories. The performance differences in novel tool generalization can be attributed primarily to training data exposure rather than architectural limitations. Our visual encoder was trained exclusively on the 115 tool categories in ToolNet, providing limited exposure to the diverse visual patterns present in novel tools. In contrast, large multimodal models like Gemini-2.0-Pro were trained on massive datasets containing millions of images across countless tool categories, enabling robust recognition of previously unseen tools.
These results highlight both the promise and limitations of our attribute-based approach for tool generalization, suggesting that while the framework can leverage learned attribute relationships to reason about novel tools, expanding the diversity of training tools and refining the attribute space could further improve generalization performance.